# Leveraging Anthropometric Measurements to Improve Human Mesh Estimation and Ensure Consistent Body Shapes

### v1

@@ 5.2 Human Mesh Estmation with Sequence Based 3D HPE and A2B Results
All evaluated HME models are working image-wise. In contrast, SOTA 3D HPE models take a long sequence of 2D poses as an input, which helps to capture movements precisely. The models are called uplifting models, since they lift 2D pose sequences to 3D pose sequences. We use the efficent SOTA 3D HPE model uplift and upsample (UU) (Einfalt et al., 2023) to estimate the 3D poses on videos. To estimate the required 2D poses from the video frames, we use ViTPose (Xu et al., 2022), a SOTA 2D pose estimation model. It is important to note that UU operates on pose sequences instead of single frames like the HME models in Section 5.1 and can leverage the information of neighboring frames to estimate a more sophisticated pose. Since we have GT 3D joints available, we can finetune the models (ViTpose for 2D HPE and UU for 3D HPE) on our data. This is also necessary to adapt the model to the dataset specific joint definitions since many 3D HPE models like UU are pretrained on datasets like Human3.6m (Ionescu et al., 2014), but those joint definitions do not match ASPset nor fit3D. We finetune both 2D and 3D HPE models on the training subsets. On the test subsets, UU achieves an MPJPE of 63.85 mm on ASPset and an MPJPE of 29.60 mm on fit3D, which is better than the best existing HME model for both datasets (see Section 5.1). However, UU only outputs 3D joints, no meshes. Moreover, a stick-figure 3D pose is not sufficient to model the pose parameters $\theta$ of the SMPL-X mesh, since some rotations are missing. Hence, it is impossible to calculate the necessary rotation parameters directly from the UU result.
@@ 5.2.1 Inverse Kinematics for Full Pose Estimation
Therefore, we use the well established approach of inverse kinematics (IK) to obtain the missing rotations by fitting an SMPL-X mesh to the 3D joint locations estimated by UU. We use the inverse kinematics approach with the VPoser extension (Pavlakos et al., 2019). VPoser is a learned prior for human poses, since the raw SMPL-X model definition allows impossible poses for humans. VPoser learned plausible poses from the large AMASS (Mahmood et al., 2019) dataset and helps IK to generate only plausible poses. IK learns the best SMPL-X parameters ($\beta$ and $\theta$) that fit the mesh to the given 3D joint locations by minimizing the error between the given joint locations and the regressed joint locations from the mesh. IK is an iterative algorithm and adjusts the pose and the shape parameters with a gradient descent minimization approach in each step. Besides the already described joint error, IK further penalizes abnormal poses with a VPoser error and extreme body shapes with a $\beta$ parameter error. We execute IK per frame, which results in a slight jitter in between the frames, but leads to more accurate joint positions. Since IK needs multiple iterations to adjust the standard T-pose parameters to achieve a pose that is roughly close to the desired UU pose, we speed up the process by initializing the pose and shape parameters with the result from the previous frame if available. This also enhances the final result slightly.
@@ 5.2.2 Experiments and Results
Figure 3: Overview of our approach. The pose and shape parameters are obtained either from IK applied to UU results (Sec. 5.2.2) or from an HME model (Sec. 5.1). In real applications, the anthropometric measurements will be taken directly from the humans . For our evaluations, we use the GT shape parameters and further experiment with the shape parameters of the respective model (IK or HME).
We evaluate different experiments in Table 5. For comparison, we mention the UU 3D HPE performance (first rows for each dataset in Tab. 5). These results correspond to stick-figure poses and not the required human meshes. Therefore, they are not directly comparable to the other results.
Our main approach is displayed in the second rows in Table 5, respectively. We evaluate the results of IK applied to the UU joint locations with original, median and pseudo GT based A2B $\beta$ parameters. The final real-world scenario corresponds to the following approach. GT anthropometric values can be measured from the athlete directly and the 3D pose can be estimated with UU and IK. The $\beta$ parameters are estimated with the A2B models. A visualization of this pipeline can be found in Figure 3. We include the best result from existing HME models in the respective last rows for comparison and provide qualitative results in Figure 1. Our approach outperforms the best existing HME model for both datasets by a large margin.
We analyze the results of the building blocks of our model in detail. Applying IK to the UU results deteriorates the UU results by nearly 4 mm for ASPset and 8 mm for fit3D (see Tab. 5, first and second rows, column orig.), but the UU result is only a stick-figure pose and not a human mesh, hence it is not sufficient for our purpose. Moreover, these results are still better than the best existing HME model.
Next, we replace the inconsistent $\beta$ parameters with the results from our A2B models. For ASPSet, using pseudo GT anthropometric measurements results in a large improvement of over 12 mm. Remarkably, this result surpasses even the original UU result by 8 mm. It seems that incorporating a clearly defined mesh helps to fix some typical errors of UU and enhance its result in case of ASPset. In general, the error on fit3D is much lower for UU based approaches. The reason might be that it consists of much more data, such that we can finetune UU for a longer time. Further, the videos are recorded in a lab in comparison to the in-the-wild videos of ASPset. Therefore, the results of ASPset are more relevant for future applications of our approach, where we assume only a few available 3D annotations and in-the-wild recordings. For fit3D, applying the A2B body shapes from pseudo GT anthropometric measurements leads to a slight decrease in performance of 0.6mm. Inconsistent shapes in the GT (see Section 3) are likely to cause this behavior. Still, our approach using a 3D HPE model and IK outperforms all existing HME models, no matter if the original inconsistent or the consistent body shapes from A2B are used.
Table 5: MPJPE results in mm on the test splits of ASPset (top) and fit3D (bottom) of our approach compared to the respective best HME model. The pose column indicates the origin of the pose. The orig column contains the result as it is estimated from the method indicated in the pose column (with inconsistent body shapes). The right block contains the results with the originally estimated $\beta$ parameters replaced by consistent ones. The measurements column indicates which anthropometric measurements are used for the A2B computation (which $beta$ parameters are used for the median computation) whose results are the replacement $\beta$ parameters in the last five columns. We highlight the best results for estimated meshes with consistent shapes.
@@
DS pose orig. measurements NN m SVR m NN n SVR n median
ASPset UU 63.85 no mesh
IK-UU 67.54 GT 56.44 56.56 55.14 55.19 -
IK-UU 67.54 IK-UU 66.92 66.60 67.25 67.12 67.16
SMPLer-X 86.02 GT 85.89 85.69 86.03 85.99 86.04
fit3D UU 29.60 no mesh
IK-UU 36.89 GT 39.36 38.69 37.46 37.89 -
IK-UU 36.89 IK-UU 40.33 38.82 38.14 38.10 38.29
Multi-HMR 72.87 GT 72.62 72.37 72.06 71.92 72.08
@@
As already described in Section 5.1, we further evaluate the capabilities of a consistent shape without available GT anthropometric measurements. The naive approach is to use the median of the estimated inconcistent $\beta$ parameters (see Tab. 5, column median). Another approach is to use the meshes created by IK applied to the UU results, compute the anthropometric measurements with B2A, calculate the median anthropometric values and convert them to $\beta$ parameters via the A2B models. Results are displayed in Table 5, row three. For ASPset, using fixed body shape parameters from A2B models based on the measurements from UU results achieves a slightly better score than the results with inconsistent body shapes. For fit3d, the MPJPE increases by another 0.6 mm, but the A2B model results are a slighly better alternative for consistent body shapes compared to the median $\beta$ parameters.
@@
Further, our approach with IK can be used to generate pseudo GT meshes for datasets with only 3D keypoint annotations. We can use these pseudo GT meshes to finetune HME models and increase their performance regarding the estimated keypoints for the specific dataset. However, these results are still worse than the results of our approach. We present the results in the supplementary.
@@ 6 Conclusion
In this paper, we address the problem of inconsistent estimated body shapes of humans in videos. We analyze the GT data of 3D pose and mesh datasets and find inconsistencies already in their annotations. Then, we propose a family of learned A2B models to convert 36 anthropometric measurements to SMPL-X $\beta$ parameters. This can be used to measure a human once and use the resulting shape of the A2B model for all evaluations. With this strategy, the body shape is accurate and consistent per person. Evaluations show that using IK on the results of a SOTA 3D HPE model to estimate the mesh pose combined with our A2B model’s shape parameters leads to superior and consistent results compared to existing HME models. Moreover, HME models also benefit from our approach. Replacing their estimated shape parameters with the A2B shape parameters created from (pseudo) GT anthropometric measurements leads to an improvement of their score and consistent body shapes.
Appendix A Anthropometric Measurements
The selection of the anthropometric measurements is mainly adopted from AnthroNet (Picetti et al., 2023). In total, 36 measurements are selected, which can be divided into 23 lengths and 13 circumferences. All measurements are taken based on the standard SMPL-X T-pose. The reference landmarks are chosen by matching the vertices on the default mesh with the landmarks defined by the anthropometric survey of the U.S. army personnel (Gordon et al., 2014). A visualization of the landmarks can be found in Figure 4 and 5(a). The lengths are calculated by computing the Euclidean distance between two landmarks, or the difference along the coordinate axis pointing upwards for certain heights. The lenghts are visualized in Figure 5(b) and 6. Table 6 lists the enclosing landmarks for each length. To measure the circumferences, we adopt the code from (Bojanic, 2023). For each measurement, a plane is created, the intersection between the mesh and the plane are extracted and the convex hull of the result is calculated. During this process, the mesh is restricted to the body part to be measured. A visualization of the circumferences can be found in Figure 7 and a list of the landmarks and the normal vectors spanning the plane in Table 7.
@@ Appendix A Anthropometric Measurements
Table 6: Definitions of lengths by their two enclosing landmarks.
@@ 7 Height Head top Center between heels
(a) Visualization of a subset of the used landmarks.
(b) Visualization of used lengths with a standard T-pose SMPL-X mesh.
Figure 5: Sideview visualizations of landmarks (5(a)) and lengths (5(b)).
Idx Circumference Normal Vector Position
1 Waist Up Belly button
2 Chest Up Nipple
3 Hip Up Pubic bone
4 Head Up Head temple
5 Neck Spine to head Adam’s apple
6/7 Upper Arm Shoulder to elbow Center of the bicep
8/9 Forearm Elbow to wrist Widest point of the forearm
10/11 Thigh Up Center of the thigh
12/13 Calf Up Widest point of the calf
Appendix B 3D Human Shape Ground Truth Analysis
We further analyze the GT shape consistency for the common datasets Human3.6m (Ionescu et al., 2014) and MPI-INF-3DHP (Mehta et al., 2017). We find that for Human3.6m, the bone lengths derived from the 3D annotations are fixed, but not for MPI-INF-3DHP. Therefore, we do not report the deviations of 3D joint annotations for Human3.6m, since there are none. We further evaluate the SMPL-X annotations for both datasets provided by NeuralAnnot (Moon et al., 2022b) which are used by HME models as GT for training. See Tables 8, 9 for details.
Table 8: GT data analysis for MPI-INF-3DHP (Mehta et al., 2017). Bone length analysis based on the 3D joint locations (left) and on SMPL-X annotations by NeuralAnnot (right). Standard deviation $\sigma$, relative standard deviation $\frac{\sigma}{\textit{avg}}$ and relative range $\frac{\max-\min}{\textit{avg}}$ of anthropometric measurements are reported. Standard deviations are given in cm, despite for the $\beta$ parameters. The values are averaged between left and right body parts and between all persons in of each dataset. The $\beta$ parameter standard deviation is averaged over all $\beta$ parameters.
3D joint annotations SMPL-X annotations
Measure $\sigma$ r. $\sigma$ r. range Measure $\sigma$ r. $\sigma$ r. range
head 0.19 1.03% 2.08% head 0.21 0.75% 4.87%
hip width 0.22 0.89% 1.80% hip circ. 1.16 1.16% 9.13 %
forearm 0.21 0.87% 1.77% forearm 0.45 1.80% 9.75%
upper arm 0.29 0.90% 1.82% arm 0.83 1.59% 8.19%
lower leg 0.60 1.49% 3.06% lower leg 1.05 2.56% 11.54%
thigh 3.83 7.91 % 41.90% thigh 0.77 2.02% 9.47%
height 2.76 1.56% 8.24%
$\beta$ param. 0.18
Table 9: GT data analysis for Human3.6m (Mehta et al., 2017): Analysis of SMPL-X annotations by NeuralAnnot. Standard deviation $\sigma$, relative standard deviation $\frac{\sigma}{\textit{avg}}$ and relative range $\frac{\max-\min}{\textit{avg}}$ of anthropometric measurements are reported. Standard deviations are given in cm, despite for the $\beta$ parameters. The values are averaged between left and right body parts and between all persons in of each dataset. The $\beta$ parameter standard deviation is averaged over all $\beta$ parameters.
SMPL-X annotations
Measure $\sigma$ r. $\sigma$ r. range
head 0.41 1.51% 10.28%
hip circ. 1.24 1.19% 8.90%
forearm 0.83 3.30% 27.93%
arm 0.77 2.58% 22.88%
lower leg 0.43 1.18% 12.20%
thigh 0.66 1.27% 9.43%
height 3.40 2.06% 15.66%
$\beta$ param. 0.20
Appendix C Evaluating A2B Models

### v3

@@ 5.2 HME with Sequence Based 3D HPE and A2B
All evaluated HME models are working image-wise. In contrast, SOTA 3D HPE models take a long sequence of 2D poses as an input, which helps to capture movements precisely. The models are called uplifting models, since they lift 2D pose sequences to 3D pose sequences. We use the efficent SOTA 3D HPE model uplift and upsample (UU) [10] to estimate the 3D poses on videos. To estimate the required 2D poses from the video frames, we use ViTPose [35], a SOTA 2D pose estimation model. It is important to note that UU operates on pose sequences instead of single frames like the HME models in Section 5.1 and can leverage the information of neighboring frames to estimate a more sophisticated pose. Since we have GT 3D joints available, we can finetune the models (ViTpose for 2D HPE and UU for 3D HPE) on our data. This is also necessary to adapt the model to the dataset specific joint definitions since many 3D HPE models like UU are pretrained on datasets like Human3.6m [16], but those joint definitions do not match ASPset nor fit3D. We finetune both 2D and 3D HPE models on the training subsets. On the test subsets, UU achieves an MPJPE of 63.85 mm on ASPset and an MPJPE of 29.60 mm on fit3D, which is better than the best existing HME model for both datasets (see Section 5.1). However, UU only outputs 3D joints, no meshes. Moreover, a stick-figure 3D pose is not sufficient to model the pose parameters $\theta$ of the SMPL-X mesh, since some rotations are missing. Hence, it is impossible to calculate the necessary rotation parameters directly from the UU result.
@@ 5.2.1 Inverse Kinematics for Full Pose Estimation
Therefore, we use the well established approach of inverse kinematics (IK) with a pose prior to obtain the missing rotations by fitting an SMPL-X mesh to the 3D joint locations estimated by UU. IK outputs the best SMPL-X parameters ($\beta$ and $\theta$) that fit the mesh to the given 3D joint locations. Details can be found in the supplementary.
@@ 5.2.2 Experiments and Results
We evaluate different experiments in Table 5. For comparison, we mention the UU 3D HPE performance (first rows for each dataset in Tab. 5). These results correspond to stick-figure poses and not the required meshes. Therefore, they are not directly comparable to the other results.
Our main approach is shown in the second rows in Table 5. We evaluate the results of IK applied to the UU joint locations with original, median, and pseudo GT based A2B $\beta$ parameters. The real-world scenario corresponds to the following. GT measurements can be measured from the athlete directly and the 3D pose can be estimated with UU and IK. The $\beta$ parameters are estimated with the A2B models. A visualization of this pipeline can be found in Figure 3. We include the best result(s) from existing HME models in the respective last rows for comparison and provide qualitative results in Figure 1. Our approach outperforms the best existing HME model for both datasets by a large margin. But we need to mention that our approach is not directly comparable to the original results of the HME models, since our approach needs the additional information of measurements, but they already exist in our scenario. However, our approach still outperforms the existing HME models even when they use the same measurments and A2B results as our model does. Further, our model provides results for all frames, which is not the case for the other HME models.
@@
DS pose orig. $\sigma$ measure NN g SVR g NN n SVR n median $\sigma$ no r. $\downarrow$
ASPset UU 63.9 no mesh
IK-UU 67.5 3.0 GT 56.4 56.6 55.1 55.2 - 0.0 0.0%
IK-UU 67.5 3.0 IK-UU 66.9 66.6 67.3 67.1 67.2 0.0 0.0%
SMPLer-X 86.0 2.9 GT 85.9 85.7 86.0 86.0 - 0.0 0.11%
fit3D UU 34.3 - no mesh
IK-UU 38.5 / 46.3 8.7 GT 41.2 / 47.5 41.3 / 46.9 38.8 / 45.3 38.7 / 45.3 - 0.0 0.0%
IK-UU 38.5 / 46.3 8.7 IK-UU 42.6 / 51.2 41.6 / 48.6 39.8 / 47.9 39.8 / 47.8 39.9 / 47.8 0.0 0.0%
Multi-HMR 74.6 / 76.1 3.3 GT 69.6 / 67.8 69.6 / 67.6 68.0 / 67.9 68.4 / 68.8 - 0.0 1.54%
Multi-HMR 74.6 / 76.1 3.3 Multi-HMR 75.8 / 77.3 76.1 / 76.7 73.9 / 75.6 74.1 / 75.8 73.9 / 75.5 0.0 1.54%
OSX 94.2 / 89.0 3.9 GT 88.9 / 83.4 88.6 / 82.3 87.1 / 81.2 87.2 / 81.1 - 0.0 3.45%
OSX 94.2 / 89.0 3.9 OSX 95.0 / 91.7 94.8 / 90.2 93.2 / 87.8 93.0 / 87.6 93.0 / 87.6 0.0 3.45%
Table 5: MPJPE and MVE results in mm on the test splits of ASPset (top) and fit3D (bottom) of our approach compared to the respective best HME model(s). For fit3D, we calculate the MVE, since we have GT meshes available. We display it as the second value in every column. The pose column indicates the origin of the pose. The orig column contains the result as it is estimated from the method indicated in the pose column (with inconsistent body shapes). The right block contains the results with the originally estimated $\beta$ parameters replaced by consistent ones. The measurements column indicates which anthropometric measurements are used for the A2B computation (which $\beta$ parameters are used for the median computation) whose results are the replacement $\beta$ parameters in the last five columns. We highlight the overall best results for estimated meshes with consistent shapes in bold and underline the best (MPJPE and MVE) results in each line. We further add the mean standard deviation of the body height and the percentage of frames with no result as in Table 3.
We analyze the results of the building blocks of our model in detail. Applying IK to the UU results deteriorates the UU results by nearly 4 mm for ASPset and 5 mm for fit3D (see Tab. 5, first and second rows, column orig.), but this step is necessary since the UU result is only a stick-figure pose and not sufficient for our purpose. Moreover, these results are still better than the best existing HME model (see last rows in Tab. 5).
Next, we replace the inconsistent $\beta$ parameters with the results from our A2B models. This is especially helpful for our approach since IK produces body shapes with high inconsistencies, as shown by the larger standard deviation of the body height compared to other HME models. For ASPSet, using pseudo GT AMs results in a large improvement of over 12 mm. Remarkably, this result surpasses even the original UU result by 8 mm. It seems that incorporating a clearly defined mesh helps to fix some typical errors of UU and enhance its result in case of ASPset. In general, the error on fit3D is much lower for UU based approaches. The reason might be that it consists of much more data, such that we can finetune UU for a longer time. Further, the videos are recorded in a lab in comparison to the in-the-wild videos of ASPset. The lab environment is very similar to the Human3.6m dataset [16], which serves as a training dataset for most recent HME models. Therefore, the results of ASPset are more relevant for future applications of our approach, where we assume only a few available 3D annotations and in-the-wild recordings. For fit3D, applying the A2B body shapes from pseudo GT AMs leads to a slight decrease in performance of 0.2 mm. Inconsistent shapes in the GT (see Section 3) are likely to cause this behavior. Still, our approach using a 3D HPE model and IK outperforms all existing HME models, no matter if the original inconsistent or the consistent body shapes from A2B are used.
Regarding the gendered meshes, we observe that the performance is slightly better for male than for female subjects. fit3D consists of two female and six male subjects. The best score of 40.2 mm for the male subjects is achieved with the SVR model. For the female subjects, the best score is 42.1 mm with the NN model.
@@
As described in Section 5.1, we further evaluate the capabilities of a consistent shape without available GT AMs. The naive approach is to use the median of the estimated inconsistent $\beta$ parameters (Tab. 5, column median). Another approach is to use the meshes created by IK applied to the UU results, compute the AMs with B2A, calculate the median AMs and convert them to $\beta$ parameters via the A2B models. Results are displayed in Table 5, row three. For ASPset, using fixed body shape parameters from A2B models based on the measurements from UU results achieves a slightly better score than the results with inconsistent body shapes. For fit3d, the MPJPE increases by 0.9 mm, but the A2B model results are a slighly better alternative for consistent body shapes compared to the median $\beta$ parameters.
@@
Further, our approach can be used to generate pseudo GT meshes for datasets with only 3D keypoint annotations. We can use these pseudo GT meshes to finetune HME models and increase their performance regarding the estimated keypoints for the specific dataset. However, these results are still worse than the results of our approach. We present the results in the supplementary.
@@ 6 Conclusion
We address the problem of inconsistent estimated basic body shapes of humans in videos. We analyze the GT data of 3D pose and mesh datasets and find inconsistencies in their annotations. Then, we propose a family of learned A2B models to convert 36 anthropometric measurements to SMPL-X $\beta$ parameters. This can be used to measure a human once (as it is established practice for athletes, our main focus) and use the resulting shape of the A2B model for all evaluations. With this strategy, the body shape is accurate and consistent per person. Evaluations show that using IK on the results of a SOTA 3D HPE model to estimate the mesh pose combined with our A2B model’s shape parameters leads to superior and consistent results compared to existing HME models. Moreover, HME models also benefit from our approach. Replacing their estimated shape parameters with the A2B shape parameters leads to an improvement of their score and consistent body shapes. However, our approach based on 3D HPE still outperforms these scores.
Supplementary Material
7 Anthropometric Measurements
@@
The selection of the anthropometric measurements is mainly adopted from AnthroNet [28]. In total, 36 measurements are selected, which can be divided into 23 lengths and 13 circumferences. All measurements are taken based on the standard SMPL-X T-pose. The reference landmarks are chosen by matching the vertices on the default mesh with the landmarks defined by the anthropometric survey of the U.S. army personnel [14]. A visualization of the landmarks can be found in Figure 4 and 5. The lengths are calculated by computing the Euclidean distance between two landmarks or the difference along the coordinate axis pointing upwards for certain heights.
Figure 5: Visualization of a subset of the used landmarks in side view.
The lenghts are visualized in Figure 6 and 7. Table 9 lists the enclosing landmarks for each length. To measure the circumferences, we adopt the code from [2]. For each measurement, a plane is created, the intersection between the mesh and the plane are extracted and the convex hull of the result is calculated. During this process, the mesh is restricted to the body part to be measured. A visualization of the circumferences can be found in Figure 8 and a list of the landmarks and the normal vectors spanning the plane in Table 6.
Figure 6: Visualization of used lengths with a standard T-pose SMPL-X mesh in side view.
Idx Circumference Normal Vector Position
1 Waist Up Belly button
2 Chest Up Nipple
3 Hip Up Pubic bone
4 Head Up Head temple
5 Neck Spine to head Adam’s apple
6/7 Upper Arm Shoulder to elbow Center of the bicep
8/9 Forearm Elbow to wrist Widest point of the forearm
10/11 Thigh Up Center of the thigh
12/13 Calf Up Widest point of the calf
8 3D Human Shape Ground Truth Analysis
We further analyze the GT shape consistency for the common datasets Human3.6m [16] and MPI-INF-3DHP [22]. We find that for Human3.6m, the bone lengths derived from the 3D annotations are fixed, but not for MPI-INF-3DHP. Therefore, we do not report the deviations of 3D joint annotations for Human3.6m, since there are none. We further evaluate the SMPL-X annotations for both datasets provided by NeuralAnnot [24] which are used by HME models as GT for training. See Tables 7, 8 for details.
3D joint annotations SMPL-X annotations
Measure $\sigma$ r. $\sigma$ r. range Measure $\sigma$ r. $\sigma$ r. range
head 0.19 1.03% 2.08% head 0.21 0.75% 4.87%
hip width 0.22 0.89% 1.80% hip circ. 1.16 1.16% 9.13 %
forearm 0.21 0.87% 1.77% forearm 0.45 1.80% 9.75%
upper arm 0.29 0.90% 1.82% arm 0.83 1.59% 8.19%
lower leg 0.60 1.49% 3.06% lower leg 1.05 2.56% 11.54%
thigh 3.83 7.91 % 41.90% thigh 0.77 2.02% 9.47%
height 2.76 1.56% 8.24%
$\beta$ param. 0.18
Table 7: GT data analysis for MPI-INF-3DHP [22]. Bone length analysis based on the 3D joint locations (left) and on SMPL-X annotations by NeuralAnnot (right). Standard deviation $\sigma$, relative standard deviation $\frac{\sigma}{\textit{avg}}$ and relative range $\frac{\max-\min}{\textit{avg}}$ of anthropometric measurements are reported. Standard deviations are given in cm, despite for the $\beta$ parameters. The values are averaged between left and right body parts and between all persons in of each dataset. The $\beta$ parameter standard deviation is averaged over all $\beta$ parameters.
SMPL-X annotations
Measure $\sigma$ r. $\sigma$ r. range
head 0.41 1.51% 10.28%
hip circ. 1.24 1.19% 8.90%
forearm 0.83 3.30% 27.93%
arm 0.77 2.58% 22.88%
lower leg 0.43 1.18% 12.20%
thigh 0.66 1.27% 9.43%
height 3.40 2.06% 15.66%
$\beta$ param. 0.20
Table 8: GT data analysis for Human3.6m [16]: Analysis of SMPL-X annotations by NeuralAnnot. Standard deviation $\sigma$, relative standard deviation $\frac{\sigma}{\textit{avg}}$ and relative range $\frac{\max-\min}{\textit{avg}}$ of anthropometric measurements are reported. Standard deviations are given in cm, despite for the $\beta$ parameters. The values are averaged between left and right body parts and between all persons in of each dataset. The $\beta$ parameter standard deviation is averaged over all $\beta$ parameters.
@@ 7 Height Head top Center between heels
Table 9: Definitions of lengths by their two enclosing landmarks.
9 Evaluating A2B Models